# Monadic Second Order Logic of the Real Order with the Set Quantifier ranging over the Borel Sets

### v3

@@
Abstract.
A celebrated 1969 theorem of Michael Rabin is that the monadic second order (MSO) theory of the real order where the monadic quantifier is allowed only to range over the closed or $F_{\sigma}$-sets, is decidable. In 1975 Saharon Shelah [20, Conjecture 7A] conjectured that if the monadic quantifier is allowed to range over the Borel subsets of the reals, the resulting MSO theory is still decidable. We confirm this conjecture.
I thank Joan Bagaria for the invitation to visit him at the University of Barcelona in April 2025. The conversation we had gave me a crucial insight. Namely, I showed Joan the idea of well-levelled orders and he proposed that this could probably be used in descriptive set theory. I went to work on this and connected it with Rabin’s work and solving the problem of MSO on the Borel sets, about which I had been thinking about on and off for a while. The idea of writing a joint paper on this did not converge, due to Joan’s other projects. I thank Andreas Blass, Jouko Väänänen and Andrés Villaveces on very useful comments on the preprint of the paper.
@@ 1. Introduction
Monadic Second Order logic (MSO) is defined like the first order logic (FO), but with an addition of set variables and the additional atomic formulas of the form $x\in{\mathbf{A}}$ and $(\exists{\mathbf{A}})\psi({\mathbf{A}})$ for a set-quantifier-free $\psi({\mathbf{A}})$. This logic is a small fragment of the Second Order Logic (SO), yet it has proven sufficiently rich to express various useful mathematical statements. The results of this paper show, for example, that for any given $\alpha<\omega_{1}$ the Determinacy of the $\alpha$-level of the Borel hierarchy is expressible as a sentence in the MSO theory S2S of the Cantor tree with the two successor functions (Corollary 6.3). On the other hand, MSO is close enough to FO to have many properties that let us work with the logic in practice, in particular many instances of decidability. The decidability of various logics is a question stemming directly from Hilbert’s programme, most famously through Hilbert’s 10th Problem and Yuri Matijashevich’s 1970 solution to it [16]. The full SO is most often undecidable. In the case of SO arithmetic, so the SO theory of $({\mathbb{N}},+,\cdot,<,0,1)$, the undecidability was proven by Alonzo Church in [4]. In contrast, the MSO theory of various structures is decidable, including the MSO theory of $({\mathbb{N}},+,\cdot,<,0,1)$, which was famously proved decidable by J. Richard Büchi in [1]. In his proof Büchi introduced an important method, which is to associate to an MSO theory the runs of an automaton. A posteriori, one could say that the most important property of MSO is its closeness to the theory of automata. This makes MSO a logic much studied in theoretical computer science. The survey article [7] by Yuri Gurevich presents many properties of the MSO logic.
A celebrated theorem of Michael Rabin in [19, Th 2.7] vastly extended Büchi’s theorem in proving that the MSO theory of the real order where the monadic quantifier is allowed only to range over closed or $F_{\sigma}$-sets, is decidable. The limits of generalisations were shown by Saharon Shelah in an important 1975 paper [20]. Shelah proved that under the assumption of the continuum hypothesis (CH), if the monadic quantifier is allowed to range over all subsets of the reals, the resulting MSO theory is undecidable (the assumption of CH was removed by Gurevich and Shelah in [8]). Shelah conjectured [20, Conjecture 7A] that if we allow the monadic quantifier to range over the Borel subsets of the reals, the resulting theory is decidable.
@@
Our methods are similar and inspired by those used by Rabin in the proof of [19, Th 2.7] to code closed and $F_{\sigma}$-sets in S2S, except that we have developed an additional coding technique. It has allowed us to code in S2S Borel sets of arbitrary complexity. That technique is a specific instance of a more general idea of well-levelled orders that this paper introduces.
The paper is organised as follows. In §2 we define well-levelled orders. They are a class of well-founded partial orders and every well-order and every tree is well-levelled. We present the basic properties of well-levelled partial orders and show some examples, to the extent that will be useful for the main result and slightly more, as a reference. In §3 we show how to use these ideas to code Borel subsets of the reals. In §4 we show how to use this coding, combined with the ideas from the original proof by Rabin, for the main decidability result. The main results are presented in §5 in Theorem 5.1 and Corollary 5.3. For the stronger version of the main result, which is Corollary 5.3, we had to leave S2S and use an additional result of Büchi [3], stating that the MSO theory of $(\omega_{1},<)$ is decidable.
@@ 2. Well-levelled orders and their basic properties
(5) The following is a main operation for the purposes of this paper. It will be generalised in Definition 2.8.
@@
By Observation 2.6(3), it suffices to show that each of the parts of the union is well-levelled. The first part is isomorphic to $^{T^{\beta}}_{y\upharpoonright[1,\beta)}$, which is well-levelled since $T^{\beta}$ is. The second part is isomorphic to $T^{\beta}\cdot<^{T}_{y(\beta)}$. We have that $^{T}_{y(\beta)}$ is a finite linear order, say of the finite size $n$. Therefore $T^{\beta}\cdot<^{T}_{y(\beta)}$ is isomorphic to $T^{\beta}\cdot n$. The conclusion follows from Observation 2.6(5) where the “tree” in question is just the linear order with $n$ elements and $P$ is $T^{\beta}$.
@@ 2.1. The Polish Topology of $T^{1+\alpha}]$
For countable ordinals $\alpha$, each $T^{1+\alpha}]$ has a naturally defined complete metric, so the resulting topology, which we shall denote by $\tau_{\alpha}$ makes $T^{1+\alpha}]$ into a Polish space. To start the proof of this, recall ([9, pg. 7]) that the Cantor space $[T]$ has a complete metric given by $d_{0}=d$ where,
@@
This metric is bounded by 1. The following classical theorem shows that this metric can be used to make any countable product $\prod_{1+\alpha}[T]$ into a Polish space. It is proved in [5, Th. 4.3.12].
@@
There is more than one way to define a compatible metric for a given Polish topology. For example [9, pg. 3] gives a different formula for the metric of a countable product of completely metrisable spaces. We like the one given in the statement of Theorem 2.13 because when applied to the countable product of the 0-1 metrics on $2=\{0,1\}$, it gives the natural metric on $2^{\omega}$.
@@ 3. Coding of the Borel Sets
Recall that $T$ stands for the Cantor tree $2^{<\omega}$. To every Borel set $H\subseteq 2^{\omega}$ we shall associate the code $c(H)$ which will be a set of branches of a structure contained in $\bigcup_{0<\alpha<\omega_{1}}T^{\alpha}$. Recall that the Borel sets are the members of the hierarchy of sets defined by the increasing sets of levels $\Sigma_{\alpha}\cup\Pi_{\alpha}$ for $\alpha<\omega_{1}$ (see [9, §11.B]). In Definition 3.3 we use the induction on $\alpha<\omega_{1}$ to describe the coding of the sets $H$ in $\Sigma_{\alpha}\cup\Pi_{\alpha}$, which will be done using sets of branches in $T^{1+\alpha}$.
@@
A branch $\rho$ in the Cantor tree $T=2^{\omega}$ is identified with the element $\bigcup\rho$ of the Cantor space $2^{\omega}$, and it is usual to denote both of these by $\rho$ and talk about them interchangeably. Then it is well known and easily verified on the basis of the definitions given so far, that any non-empty closed set $H\subseteq 2^{\omega}$ is exactly (identified with) the set $[S]$ of branches of a subtree $S$ of $T$ of height $\omega$, and vice versa, any such set of branches forms a closed set. See [9, §2.B] for a detailed exposition. We use the notation $T_{H}$ for a tree $S$ obtained from the closed set $H$ in this manner. In other words, $T_{H}=\{s\in T:\,[s]\cap H\neq\emptyset\}$.
We then may say that open sets are the sets of the form $2^{\omega}\setminus[T_{H}]$ for some $H$ closed. We therefore obtain that both open and closed sets are identified with, which we will name “coded by”, exactly themselves.
@@
It remains to deal with the case $H_{0}\cap H_{1}=\emptyset$. This implies that there is $n_{0}$ such that for all $n\geq n_{0}$ we have that $H^{0}_{n}\cap H^{1}_{n}=\emptyset$ and hence by the induction hypothesis we can assume that $c(H^{0}_{n_{0}})\cap(H^{1}_{n_{0}})=\emptyset$. By the construction of codes at limit ordinals it follows that $c(H_{0})\cap(H_{1})=\emptyset$.
@@ 4.1. The recursive correspondance between MSO formulas over $T$ and Rabin automata
It was shown in [19, pg.8] that both $\leq$ and $\leq_{\rm lex}$ are MSO-definable using just $r_{0}$ and $r_{1}$. It follows that the MSO theory of the ordered Cantor tree along with the lexicographic order is a sub-theory of $S2S$. The main theorem of [19] is Theorem 1.7, which shows that to every formula of $S2S$ there is an associated Rabin automaton which represents the formula. To recall, this means that there is an elementary recursive procedure which to every such formula $F({\bf A}_{1},\ldots{\bf A}_{n})$, where ${\bf A}_{1},\ldots{\bf A}_{n}$ are set variables, associates a Rabin automaton ${\mathfrak{A}}_{F}$ which accepts exactly the set of all representations of tuples $(A_{1},\ldots A_{n})$ such that ${\mathfrak{N}}_{2}\vDash F(A_{1},\ldots A_{n})$. Let us recall what is meant by the Rabin automaton and representations (see [19, pg. 6]).
An automaton $\mathfrak{A}$} is an object consisting of finitely many states forming the set $S$ (its “hardware”) of which some are designated as initial states, forming the set $S_{0}$, and some are designated as final stages, forming the set $F^{*}$. Rabin automata have an associated set of labels, denoted $L$, which for the example of $n$-place formulas as above will be the set $2^{n}$. The “software” of the automaton or its “programme” is a table of moves, which is a function $M$ which to every pair consisting of a state and a label, associates a subset of $S\times S$. Runs of a Rabin automaton are defined over trees $T_{s}=\{t\in T:\,s\leq t\}$ (so the basic open sets in $T$) labelled by $L$, so pairs of the form $(T_{s},v)$ where $v$ is some function $v:T_{s}\to L$.
@@ 4.3. Translating into $S2S$
The proof is inspired by Rabin’s proof of [19, Th 1.12] mentioned above.
Proof.
@@
Case $\alpha>0$ a limit ordinal. Let us first work with $\alpha=\omega$. The way that Rabin showed in [19, Th 1.12] that we can MSO-interpret $S\omega S$ within $S2S$ is by finding an MSO-definable subset $A^{*}$ of $T$ such that
@@
We would like to generalise this to any given countable limit ordinal $\alpha$, but we need a different approach since it is not obvious how to MSO-define “being of order type $\alpha$”. To see that this possible, we first recall Büchi’s result from [2] which says that for any $\beta<\omega_{1}$ to any MSO-formula $F(\bar{X})$ over the structure $(\beta,<)$, there corresponds a Büchi automaton $\mathfrak{A}$ which accepts the input $\bar{X}$ iff $F(\bar{X})$ holds, and this correspondence is also true in the other sense. Then we recall Rabin’s result in [19, §3] that to the Büchi automaton $\mathfrak{A}$ there corresponds a Rabin automaton $\mathfrak{B}$ and a recursive transformation $f$ of the input space, such that $\mathfrak{B}$ accepts the input $f(\bar{X})$ iff $\mathfrak{A}$ accepts the input $\bar{X}$. (This result uses Mc Naughton’s theorem from [17] which transforms Büchi automata to Müller ones). Finally, Rabin’s result from [19] gives that that there is an MSO-formula $F_{T}(\bar{Y})$ such that $F_{T}(\bar{Y})$ is true over $T$ iff $\mathfrak{B}$ accepts the input $(\bar{Y})$. Putting these results together, we obtain the interpretation of the MSO theory of $(\beta,<)$ in the MSO theory of $(T,\leq,\leq_{\rm lex})$.
@@
Hence $G_{\alpha}(\varrho)$ holds if $\varrho$ is an $\alpha$-sequence of elements of $T$ which is eventually $\Lambda$. It follows that there is a recursive bijection between $I_{\alpha}=\{\varrho:\,G_{\alpha}(\varrho)\}$ and $T^{\alpha}$. Hence it suffices to define $f^{\alpha}$ on $I_{\alpha}$. For every $\varrho\in I_{\alpha}$ and $\beta<\alpha$ we can MSO-define $\varrho\upharpoonright\beta$ with parameter $\varrho$, since $\beta$ is MSO-definable, “being an ordinal” is MSO-definable and “being an ordinal $\alpha$” is MSO-definable. We can also MSO-define $\beta(\varrho)$ as the first such that for all $\beta>\beta(\varrho)$ we have $\varrho(\beta)=\lambda$. Then it suffices to let $f^{\alpha}(\varrho)=f^{\beta(\varrho)}(\varrho)$. We then have

### v4

# Monadic Second Order Logic of the Real Order with the Set Quantifiers ranging over the Borel Sets

@@
Monadic Second Order Logic of the Real Order with the Set Quantifiers ranging over the Borel Sets
Mirna Džamonja base-name Logique Consult, 7 Impasse Chartière, 75005 Paris, France logique.consult@gmail.com https://www.logiqueconsult.eu
Аннотация.
[…]
I thank Joan Bagaria for the invitation to visit him at the University of Barcelona in April 2025. The conversation we had gave me a crucial insight. Namely, I showed Joan the idea of well-levelled orders and he proposed that this could probably be used in descriptive set theory. I went to work on this and connected it with Rabin’s work and solving the problem of MSO on the Borel sets, about which I had been thinking about on and off for a while. The idea of writing a joint paper on this did not converge, due to Joan’s other projects. I thank Andreas Blass, Jouko Vддnдnen and Andrйs Villaveces on very useful comments on the preprint of the paper.
@@ 1. Introduction
Monadic Second Order logic (MSO) is defined like the first order logic (FO), but with an addition of set variables and the additional atomic formulas of the form $x\in{\mathbf{A}}$ and $(\exists{\mathbf{A}})\psi({\mathbf{A}})$ for a set-quantifier-free $\psi({\mathbf{A}})$. This logic is a small fragment of the Second Order Logic (SO), yet it has proven sufficiently rich to express various useful mathematical statements. The results of this paper show, for example, that for any given $\alpha<\omega_{1}$ the Determinacy of the $\alpha$-level of the Borel hierarchy is expressible as a sentence in the MSO theory S2S of the Cantor tree with the two successor functions (Corollary 6.3). On the other hand, MSO is close enough to FO to have many properties that let us work with the logic in practice, in particular many instances of decidability. The decidability of various logics is a question stemming directly from Hilbert’s programme, most famously through Hilbert’s 10th Problem and Yuri Matijashevich’s 1970 solution to it [17]. The full SO is most often undecidable. In the case of SO arithmetic, so the SO theory of $({\mathbb{N}},+,\cdot,<,0,1)$, the undecidability was proven by Alonzo Church in [4]. In contrast, the MSO theory of various structures is decidable, including the MSO theory of $({\mathbb{N}},+,\cdot,<,0,1)$, which was famously proved decidable by J. Richard Bьchi in [1]. Bьchi uses an important method, which is to associate to an MSO theory the runs of an automaton. This method was introduced by Boris (Boaz) Trakhtenbrot in [22], and in fact, a posteriori, one could say that the most important property of MSO is its closeness to the theory of automata. This makes MSO a logic much studied in theoretical computer science. The survey article [8] by Yuri Gurevich presents many properties of the MSO logic.
A celebrated theorem of Michael Rabin in [20, Th 2.7] vastly extended Bьchi’s theorem in proving that the MSO theory of the real order where the monadic quantifier is allowed only to range over closed or $F_{\sigma}$-sets, is decidable. The limits of generalisations were shown by Saharon Shelah in an important 1975 paper [21]. Shelah proved that under the assumption of the continuum hypothesis (CH), if the monadic quantifier is allowed to range over all subsets of the reals, the resulting MSO theory is undecidable (the assumption of CH was removed by Gurevich and Shelah in [9]). Shelah conjectured [21, Conjecture 7A] that if we allow the monadic quantifier to range over the Borel subsets of the reals, the resulting theory is decidable.
@@
Our methods are similar and inspired by those used by Rabin in the proof of [20, Th 2.7] to code closed and $F_{\sigma}$-sets in S2S, except that we have developed an additional coding technique. It has allowed us to code in S2S Borel sets of arbitrary complexity. That technique is a specific instance of a more general idea of well-levelled orders that this paper introduces.
The paper is organised as follows. In §2 we define well-levelled orders. They are a class of well-founded partial orders and every well-order and every tree is well-levelled. We present the basic properties of well-levelled partial orders and show some examples, to the extent that will be useful for the main result and slightly more, as a reference. In §3 we show how to use these ideas to code Borel subsets of the reals. In §4 we show how to use this coding, combined with the ideas from the original proof by Rabin, for the main decidability result. The main results are presented in §5 in Theorem 5.1 and Corollary 5.3. For the stronger version of the main result, which is Corollary 5.3, we had to leave S2S and use an additional result of Bьchi [3], stating that the MSO theory of $(\omega_{1},<)$ is decidable.
@@ 2. Well-levelled orders and their basic properties
(5) The following is a main operation for the purposes of this paper. It will be generalised in Definition 2.8. The author used this construction in [5] to solve a question in combinatorics (posed by R. Ahlswede, Peter Erdős and N. Graham) .
@@
By Observation 2.6(3), it suffices to show that each of the parts of the union is well-levelled. The first part is isomorphic to $^{T^{\beta}}_{y\upharpoonright[1,\beta)}$, which is well-levelled since $T^{\beta}$ is. The second part is isomorphic to $T^{\beta}\cdot<^{T}_{y(\beta)}$. We have that $^{T}_{y(\beta)}$ is a finite linear order, say of the finite size $n$. Therefore $T^{\beta}\cdot<^{T}_{y(\beta)}$ is isomorphic to $T^{\beta}\cdot n$. The conclusion follows from Observation 2.6(5) where the ‘‘tree’’ in question is just the linear order with $n$ elements and $P$ is $T^{\beta}$.
@@ 2.1. The Polish Topology of $T^{1+\alpha}]$
For countable ordinals $\alpha$, each $T^{1+\alpha}]$ has a naturally defined complete metric, so the resulting topology, which we shall denote by $\tau_{\alpha}$ makes $T^{1+\alpha}]$ into a Polish space. To start the proof of this, recall ([10, pg. 7]) that the Cantor space $[T]$ has a complete metric given by $d_{0}=d$ where,
@@
This metric is bounded by 1. The following classical theorem shows that this metric can be used to make any countable product $\prod_{1+\alpha}[T]$ into a Polish space. It is proved in [6, Th. 4.3.12].
@@
There is more than one way to define a compatible metric for a given Polish topology. For example [10, pg. 3] gives a different formula for the metric of a countable product of completely metrisable spaces. We like the one given in the statement of Theorem 2.13 because when applied to the countable product of the 0-1 metrics on $2=\{0,1\}$, it gives the natural metric on $2^{\omega}$.
@@ 3. Coding of the Borel Sets
Recall that $T$ stands for the Cantor tree $2^{<\omega}$. To every Borel set $H\subseteq 2^{\omega}$ we shall associate the code $c(H)$ which will be a set of branches of a structure contained in $\bigcup_{0<\alpha<\omega_{1}}T^{\alpha}$. Recall that the Borel sets are the members of the hierarchy of sets defined by the increasing sets of levels $\Sigma_{\alpha}\cup\Pi_{\alpha}$ for $\alpha<\omega_{1}$ (see [10, §11.B]). In Definition 3.3 we use the induction on $\alpha<\omega_{1}$ to describe the coding of the sets $H$ in $\Sigma_{\alpha}\cup\Pi_{\alpha}$, which will be done using sets of branches in $T^{1+\alpha}$.
@@
A branch $\rho$ in the Cantor tree $T=2^{\omega}$ is identified with the element $\bigcup\rho$ of the Cantor space $2^{\omega}$, and it is usual to denote both of these by $\rho$ and talk about them interchangeably. Then it is well known and easily verified on the basis of the definitions given so far, that any non-empty closed set $H\subseteq 2^{\omega}$ is exactly (identified with) the set $[S]$ of branches of a subtree $S$ of $T$ of height $\omega$, and vice versa, any such set of branches forms a closed set. See [10, §2.B] for a detailed exposition. We use the notation $T_{H}$ for a tree $S$ obtained from the closed set $H$ in this manner. In other words, $T_{H}=\{s\in T:\,[s]\cap H\neq\emptyset\}$.
We then may say that open sets are the sets of the form $2^{\omega}\setminus[T_{H}]$ for some $H$ closed. We therefore obtain that both open and closed sets are identified with, which we will name ‘‘coded by’’, exactly themselves.
@@
It remains to deal with the case $H_{0}\cap H_{1}=\emptyset$. This implies that there is $n_{0}$ such that for all $n\geq n_{0}$ we have that $H^{0}_{n}\cap H^{1}_{n}=\emptyset$ and hence by the induction hypothesis we can assume that $c(H^{0}_{n_{0}})\cap c(H^{1}_{n_{0}})=\emptyset$. By the construction of codes at limit ordinals it follows that $c(H_{0})\cap c(H_{1})=\emptyset$.
@@ 4.1. The recursive correspondance between MSO formulas over $T$ and Rabin automata
It was shown in [20, pg.8] that both $\leq$ and $\leq_{\rm lex}$ are MSO-definable using just $r_{0}$ and $r_{1}$. It follows that the MSO theory of the ordered Cantor tree along with the lexicographic order is a sub-theory of $S2S$. The main theorem of [20] is Theorem 1.7, which shows that to every formula of $S2S$ there is an associated Rabin automaton which represents the formula. To recall, this means that there is an elementary recursive procedure which to every such formula $F({\bf A}_{1},\ldots{\bf A}_{n})$, where ${\bf A}_{1},\ldots{\bf A}_{n}$ are set variables, associates a Rabin automaton ${\mathfrak{A}}_{F}$ which accepts exactly the set of all representations of tuples $(A_{1},\ldots A_{n})$ such that ${\mathfrak{N}}_{2}\vDash F(A_{1},\ldots A_{n})$. Let us recall what is meant by the Rabin automaton and representations (see [20, pg. 6]).
An automaton $\mathfrak{A}$} is an object consisting of finitely many states forming the set $S$ (its ‘‘hardware’’) of which some are designated as initial states, forming the set $S_{0}$, and some are designated as final stages, forming the set $F^{*}$. Rabin automata have an associated set of labels, denoted $L$, which for the example of $n$-place formulas as above will be the set $2^{n}$. The ‘‘software’’ of the automaton or its ‘‘programme’’ is a table of moves, which is a function $M$ which to every pair consisting of a state and a label, associates a subset of $S\times S$. Runs of a Rabin automaton are defined over trees $T_{s}=\{t\in T:\,s\leq t\}$ (so the basic open sets in $T$) labelled by $L$, so pairs of the form $(T_{s},v)$ where $v$ is some function $v:T_{s}\to L$.
@@ 4.3. Translating into $S2S$
The proof is inspired by Rabin’s proof of [20, Th 1.12] mentioned above.
Доказательство.
@@
Case $\alpha>0$ a limit ordinal. Let us first work with $\alpha=\omega$. The way that Rabin showed in [20, Th 1.12] that we can MSO-interpret $S\omega S$ within $S2S$ is by finding an MSO-definable subset $A^{*}$ of $T$ such that
@@
We would like to generalise this to any given countable limit ordinal $\alpha$, but we need a different approach since it is not obvious how to MSO-define ‘‘being of order type $\alpha$’’. To see that this possible, we first recall Büchi’s result from [2] which says that for any $\beta<\omega_{1}$ to any MSO-formula $F(\bar{X})$ over the structure $(\beta,<)$, there corresponds a Büchi automaton $\mathfrak{A}$ which accepts the input $\bar{X}$ iff $F(\bar{X})$ holds, and this correspondence is also true in the other sense. Then we recall Rabin’s result in [20, §3] that to the Büchi automaton $\mathfrak{A}$ there corresponds a Rabin automaton $\mathfrak{B}$ and a recursive transformation $f$ of the input space, such that $\mathfrak{B}$ accepts the input $f(\bar{X})$ iff $\mathfrak{A}$ accepts the input $\bar{X}$. (This result uses Mc Naughton’s theorem from [18] which transforms Büchi automata to Mьller ones). Finally, Rabin’s result from [20] gives that that there is an MSO-formula $F_{T}(\bar{Y})$ such that $F_{T}(\bar{Y})$ is true over $T$ iff $\mathfrak{B}$ accepts the input $(\bar{Y})$. Putting these results together, we obtain the interpretation of the MSO theory of $(\beta,<)$ in the MSO theory of $(T,\leq,\leq_{\rm lex})$.
@@
Hence $G_{\alpha}(\varrho)$ holds if $\varrho$ is an $\alpha$-sequence of elements of $T$ which is eventually $\Lambda$. It follows that there is a recursive bijection between $I_{\alpha}=\{\varrho:\,G_{\alpha}(\varrho)\}$ and $T^{\alpha}$. Hence it suffices to define $f^{\alpha}$ on $I_{\alpha}$. For every $\varrho\in I_{\alpha}$ and $\beta<\alpha$ we can MSO-define $\varrho\upharpoonright\beta$ with parameter $\varrho$, since $\beta$ is MSO-definable, ‘‘being an ordinal’’ is MSO-definable and ‘‘being an ordinal $\alpha$’’ is MSO-definable. We can also MSO-define $\beta(\varrho)$ as the first such that for all $\beta>\beta(\varrho)$ we have $\varrho(\beta)=\lambda$. Then it suffices to let $f^{\alpha}(\varrho)=f^{\beta(\varrho)}(\varrho)$. We then have